OpenStudy (anonymous):

If the legs of a right triangle have length 6 inches and 8 inches, what is the length of the hypotenuse in inches? 11 13 10 16

OpenStudy (anonymous):

Do you know the Pythagorean theorem?

OpenStudy (anonymous):

I'm just learning it and lesson is really bad at explaining.

OpenStudy (anonymous):

\[x^2+b^2=c^2\] \[6^2+8^2=c^2\] \[36+64=c^2\] \[100=c^2\] \[\sqrt{100} = \sqrt{c^2}\] c = 10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The hypotenuse is the longest side and is always represented as C so you plug in 6 and 8 for A and B then solve for C.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you making it easier to understand then the lesson does.

OpenStudy (wwwhhhaaattt?):

@NobleWho WERE FAMILY YAYAYAYAYAY

OpenStudy (anonymous):

Yay!